OpenStudy (anonymous):

Find dy/dx for y=sec(1+x^2).

OpenStudy (anonymous):

Solution: http://screencast.com/t/NEYiosPUi

OpenStudy (loser66):

is it not the same 2 layers derivative ? (sec(u))' = u' sec u tan u = 2x sec (1+x^2) tan (1+x^2)? either way is right, right?

OpenStudy (anonymous):

Abselutely @Loser66 . I made a mistake . The derivetive of cos(1+x^2) should be -2x sin(1+x^2) in my solution.

OpenStudy (anonymous):

\[y=\sec \left( 1+x ^{2} \right)\] differentiate w.r.t. x, \[\frac{ dy }{ dx }=\sec \left( 1+x ^{2} \right)\tan \left(1 +x ^{2} \right)2x\]